Maria and Eduardo are traveling to a business conference. The trip takes 2 hr for Maria and 2.5 hr for Eduardo, since he lives 40 mi farther away. Eduardo travels 5 mph faster than Maria. Find their average rates.
Answer by ankor@dixie-net.com(15652)

Maria and Eduardo are traveling to a business conference. The trip takes 2 hr for Maria and 2.5 hr for Eduardo, since he lives 40 mi farther away. Eduardo travels 5 mph faster than Maria. Find their average rates.
:
Let s = M's speed
Then
(s+5) = E's speed
:
Write a distance equation: Dist = Time * speed
:
E's trip - M's trip = 40
2.5(s+5) - 2s = 40
:
2.5s + 12.5 - 2s = 40
:
2.5s - 2s = 40 - 12.5
:
.5s = 27.5
:
s = 27.5/.5
:
s = 55 mph is M's speed
then
55 + 5 = 60 mph is E's speed
:
:
Check our solutions by finding the distances
M: 2 * 55 = 110 mi
E: 2.5 * 60 = 150 mi which is 40 mi further
:
Did the make sense to you. Not that hard, right?